### v1

# Coherent Control of an Optical Quantum Dot Using Phonons and Photons

Abstract
Genuine quantum-mechanical effects are readily observable in modern optomechanical systems comprising bosonic (“classical”) optical resonators. Such systems have enabled laser-cooling of mesoscopic objects, generation of squeezed light, and the conversion of photons between microwave and optical modes. Here we demonstrate unique advantages of optical two-level systems, or qubits, for optomechanics. The qubit state can be coherently controlled using an immense variety of driving schemes including both phonons and resonant or detuned photons. We experimentally demonstrate this using charge-controlled InAs quantum dots (QDs) in surface-acoustic-wave resonators. Time-correlated single-photon counting measurements reveal the control of QD population dynamics using engineered optical pulses and mechanical motion. As a first example, we show how this can improve signal-to-background scattering in microwave-to-optical transduction processes. Specifically, we tailor the scheme so that mechanically assisted photon scattering is enhanced over the direct detuned photon scattering from the optical system. These differences are greatly amplified by strategic temporal pulse shaping. Quantum-mechanical calculations show good agreement with our experimental results and provide guidance for adapting these schemes to small phonon occupancies relevant for microwave-to-optical quantum transduction.
Optomechanics underlies major technological advancements ranging from single particle and atom trapping [1] to gravitational-wave detection [2, 3]. As detection sensitivities and fabrication capabilities have improved, cavity optomechanics has been extended to explore and exploit more subtle quantum-mechanical effects [2, 4], enabling laser-cooling of mesoscopic objects [5, 6, 7, 8], new sources of squeezed light [9], transduction between microwaves and optical light [10, 11, 12, 13, 14, 7, 15], and entanglement between optical and mechanical [16, 17] or optical and microwave modes [18, 19]. Notably, all such previous demonstrations have exploited bosonic (“classical”) optical resonators, for example, whispering-gallery-mode resonators [5], Fabry-Perot resonators [10, 8], and photonic-crystal nanobeams [6, 11, 12, 13, 19, 15]; their purpose is essentially to confine the optical field to the region where the optomechanical interaction occurs, or to enforce frequency-dependent optical admittances that depend sensitively on mechanical motion. For applications where single-quantum interactions are most relevant, such as quantum microwave-optical transduction, single-phonon generation [20] or single-photon generation [21], optical two-level systems (TLSs) — appropriately referred to as qubits — may be beneficial because they inherently limit optical interactions to single-photon levels. They also enable coherent control schemes that are not possible in bosonic optical resonators. Optical TLSs and related three-level systems can resonantly couple with mechanical modes, offering potential routes to efficient optomechanical cooling and phonon “lasing” [22].
Semiconductor quantum dots (QDs) are excellent realizations of optical TLSs [23, 24], exhibiting nearly gigahertz emission rates, near-unity quantum yields, near-transform-limited linewidths, high single-photon purity and high single-photon indistinguishability [25]. They have been used to generate nonclassical states of light [26], spin-photon entanglement [27, 28, 29], and photon-photon entanglement [30, 31]. Demonstrations of unconventional coherent excitation schemes using two ultrafast red-detuned laser pulses highlight the remarkable degree of control possible with QDs [21, 32]. Recent efforts involve integrating QDs into larger multifunctional hybrid quantum systems [33, 34, 35, 36]. QDs and other solid-state quantum emitters, e.g., diamond color centers, have more recently been explored as light-matter intermediaries in quantum optomechanical systems [22, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48]. Oscillations between excited orbital states in diamond color centers have been observed using multi-phonon processes and introduced as a method for dynamically mitigating decoherence [49]. To date, we know of no demonstrations of coherent optical and mechanical control of a coupled QD-phonon system. Such an effect could be used to improve signal-to-background performance in applications such as microwave-to-optical transduction by controlling the probabilities of direct-excitation scattering compared to phonon-assisted scattering from the QD. These same methods may be immediately applied to “swing-up” schemes to coherently prepare the QD’s exciton (biexciton) state for single-photon (entangled-photon) generation using a single nonresonant pump which can then be spectrally filtered [50].
Here, we use frequency- and time-tailored optical pulses, driven surface acoustic wave (SAW) cavity phonons, and time-domain optical measurements to experimentally demonstrate phonon-assisted coherent dynamics in a single QD. That is, QD excitation dynamics are controlled using both phonons and photons. Mechanical motion has large effects on TLS trajectories up to approximately 1 ns after the onset of optical driving, even after substantial radiative relaxation. As a first example — relevant to microwave-to-optical transduction — we show that optical pulse shapes and durations can be optimized such that phonon-assisted scattering is substantially enhanced compared to background scattering from direct detuned excitation. Quantum-mechanical calculations show good agreement with experiment, together indicating that large signal-to-background improvements are possible in these systems with optimized pulse shaping and device performance. There is no analogue in optomechanical systems comprising classical optical resonators.
I Visualizing phonon-assisted TLS dynamics
Current realizations of hybrid QD-mechanical systems are operated at timescales much longer than relevant dephasing times, i.e. via continuous-wave (CW) pumping [41, 42, 34, 43, 47], or with nonresonant optical pumping [51, 38, 44], and therefore these coherent properties are obscured. In a few cases, unique TLS signatures have been observed through second-order photon correlation measurements [41, 47]. Dynamics in phonon-assisted scattering have been observed, but the optical TLS coherence was not exploited or explored [42]. In the CW case, the QD merely plays a role analogous to the optical resonators in the above-mentioned bosonic systems. In fact, the two-level nature of the QD could even be considered a disadvantage in these cases, as saturation can occur under the strong optical pumping required to get significant optomechanical coupling rates. This is remedied by using short optical pulses and well-known coherent control methods, as we show here.
Our system comprises a single optical QD with transition frequency $\omega_{\text{QD}}$ subject to an optical drive (frequency $\omega_{\text{pump}}$=$\omega_{\text{QD}}$+$\Delta$ with detuning $\Delta$) in a single-mode SAW cavity (frequency $\omega_{\text{SAW}}$). The SAW strain field parametrically modulates the QD’s frequency via deformation potential coupling [52, 37, 53, 54]. This coupling is described by a rate $g$, leading to an instantaneous QD frequency $\omega_{\text{QD}}$+$g\cos(\omega_{\text{SAW}}t)$. Importantly, this phonon-induced frequency modulation leads to discernible variations in the TLS occupancy when the QD is optically driven (Fig. 1). In practice, an optimized pulse shape would be developed to prepare the TLS in a desired state, then successive processes would be triggered on photons collected after this pulse. A more detailed theoretical description of this system is provided in Methods and in the Supplementary Information.
[…]
Figure 1: (a) Top row: Bloch-sphere representation of the two-level system time evolution for: $i$) direct resonant excitation ($\Delta$=0), $ii$) direct red-detuned excitation ($\Delta$=$-\omega_{\text{SAW}}$), $iii$) phonon-assisted red-detuned excitation assuming a large coherent phonon population ($\Delta$=$-\omega_{\text{SAW}}$, $g/2\pi$=2 GHz). In all cases a constant $\Omega_{0}/2\pi$=1.5 GHz is used with the TLS initiated in its ground state at time $t$=0. Bottom row: Expected excited-state occupancy $\langle\rho_{ee}(t)\rangle$ for the same three cases. Calculations in panel a) do not include radiative relaxation or dephasing. (b) Time-dependent comparison of $\langle\rho_{ee}(t)\rangle$ between cases $ii$) (brown dashed curve) and $iii$) (red). Case $iii$) is shown for $g/2\pi$=2 GHz (pink) and $g/2\pi$=100 MHz (red) corresponding approximately to approximately $10^{3}$ phonons with $g_{0}/2\pi$=3 MHz. This calculation assumes an idealized square optical pulse shape and a radiative relaxation rate of $2\pi\times$320 MHz with no additional dephasing.
II Characterizing basic device performance
A fabricated QD-SAW optomechanical device is shown in Fig. 2a. Complete fabrication details are described in ref [46]. SAWs are confined between two acoustic mirrors forming an acoustic Fabry-Perot cavity at frequency $\omega_{\text{SAW}}/2\pi$$\approx$3.5 GHz with a quality factor $Q$$\approx$20,000 and finesse $\approx$130. The cavity both enhances the QD-SAW interaction compared to that on a bare surface [43] and allows us to parameterize the QD-SAW interaction by the single-phonon coupling rate $g_{0}$ and equilibrium phonon number $\bar{n}_{\text{SAW}}$. The SAW cavity mode is coherently populated with $\bar{n}_{\text{SAW}}$ phonons by driving an interdigital transducer (IDT) with an external microwave source.
Figure 2: (a) Optical microscope image of the device. IDT: interdigital transducer. CPW: coplanar waveguide. (b) Schematic cross-section (not to scale) of the region under the Schottky top gate (indicated by a dotted white line in panel a). A single InAs QD mediates an optomechanical interaction between SAWs ($\omega_{\text{SAW}}$) and the optical pump ($\omega_{\text{pump}}$). DBR: distributed Bragg reflector. (c) Frequency-domain illustration of the system with a pulsed pump bandwidth of 600 MHz (black curve, resonant; red curve, red detuned) and a QD linewidth of 450 MHz (gray region). Three different scattering processes are illustrated with rates $\Omega_{0}$ (resonant direct excitation), $\Omega$ (detuned direct excitation) and $\Gamma$ (detuned phonon-assisted excitation), described in the main text. (d) Two types of resonance fluorescence (‘RF’) measurements from a single QD. Top panel: Total scattered photon counts while slowly varying the CW pump detuning $\Delta$ without (brown) and with large steady-state phonon occupancies ($g/2\pi$=1.09 GHz, dark red; $g/2\pi$=1.93 GHz, lighter red). Bottom panel: Frequency spectrum of scattered photons under red-detuned ($\Delta$=$-\omega_{\text{SAW}}$) CW optical pumping. The $g$ values and corresponding colors are the same in both panels. In each panel, spectra have been normalized to the maximum counts in the dataset. Gray region: pump background. (e) Exemplary experimental time-domain optical signal (counts per time bin) under $\Delta$=0 pulsed excitation acquired with a 15-second integration time. Important experimental features are illustrated. Blue curve: QD luminescence signal with decay time $\tau_{\text{QD}}$=1 ns. Gray region; pump background, showing the pulse shape and duration $\tau_{\text{pulse}}$$\approx$800 ps.
For good QD performance, QDs are electrostatically stabilized and controlled by a vertical diode (Fig. 2b) [55, 46]. This enables tuning of the QD’s charge state and wavelength with applied gate bias voltage (Supplementary Information) [56, 57] and reduces QD absorption linewidths to $\gamma_{\text{QD}}/2\pi$$<$350 MHz in best cases [58, 59, 57]. Under CW (narrowband) pumping, direct and phonon-assisted excitation can be fairly well isolated by changing the pump frequency. Under pulsed pumping, the increased pump bandwidth makes it more difficult to isolate these processes spectrally (Fig. 2c), but it opens the new possibility of coherent population control through pulse tuning as we next demonstrate.
We interrogate a singly charged QD at 975.934 nm with an approximately 800 MHz linewidth (full width at half maximum) measured over a timescale of several seconds. This broadening beyond the radiative lifetime limit is likely due to slow (sub-MHz) spectral wandering [60, 59], as we will describe in more detail in a later section. Nonetheless, our system satisfies the condition $\omega_{\text{SAW}}$$>$$\gamma_{\text{QD}}$ — known as the resolved-sideband regime — as verified by measuring the absorption (Fig. 2d; top panel) and scattering spectra (Fig. 2d; bottom panel) of the single QD under SAW modulation. Well-defined spectral sidebands at $\Delta$=$\pm$$\omega_{\text{SAW}}$ correspond to single-phonon-assisted optical scattering from the QD [52, 61, 43]. Our main optical measurements involve exciting the QD with frequency-tuned and time-tailored optical pulses (Methods). A CW pump with variable detuning $\Delta$ originates from a narrow-linewidth ($\sim$10 kHz) laser. Pulses with durations of approximately 1 ns are generated by passing the CW pump through two electro-optic modulators and a tunable spectral filter (Methods). The time distribution of scattered photons is resolved using time-correlated single-photon counting techniques, as illustrated in Fig. 2e.
III Resolving phonon-assisted excitation dynamics
The long ($\approx$1 ns) pulse duration allows us to directly observe, within approximately 10 seconds, the QD’s average population dynamics as the QD emits photons at random times during each pulsed-excitation event. These time-domain traces of the QD’s dynamics are each composed of several million photon detection events (Methods). Series of such measurements under red-detuned ($\Delta$=$-\omega_{\text{SAW}}$) pumping with $g$ varying between approximately 0 (only thermal occupancy) and 2 GHz are shown in Figs. 3a-b. Measured count rates are converted to estimated QD occupancies using a calibration procedure described in Supplementary Information. For the remainder of this article, we will refer to the ratio of occupancies as the “enhancement”, $c_{g}(t)$=$s_{g}(t)/s_{g=0}(t)$ where $s_{g}(t)$ is the signal at time $t$ with coupling rate $g$. This enhancement measures how well the presence or absence of phonons can be detected with the optical signal.
Figure 3: (a,b) Experimental QD dynamics under red-detuned ($\Delta$=$-\omega_{\text{SAW}}$) pumping (cases $ii$ and $iii$) with two different optical pulse shapes: a) a pulse with a gradually decreasing intensity; b) a nearly square pulse with an abruptly decreasing intensity. Top row: measured pulse intensities for the two cases. Approximate peak resonant Rabi frequencies are specified in each panel. Center row: measured time-resolved QD occupancy. Trajectories for a range of QD-SAW coupling rates $g$ are shown ($g$ specified next to a subset of the curves). Bottom row: occupancy ratios (“enhancements”), $c_{g}(t)$, for each of the $g$$>$0 curves compared to the $g$=0 curve. Gray region indicates where the dynamics are dominated by free evolution (decay) of the TLS. (c,d) Numerical calculations corresponding approximately to the two experimental scenarios presented in panels a,b. Calculations assume radiative relaxation at 320 MHz and no additional dephasing. Calculations for the “exponential” pulse are truncated at 1.25 ns since oscillations would formally continue indefinitely leading to indefinitely large enhancements at long times. Note that the occupancy ratio in panel c is presented on two independently scaled axes because maxima span several orders of magnitude.
We first discuss results using a pulse with a gradually decreasing intensity (Fig. 3a). Experimentally measured pulse shapes are shown in the top panels. The QD occupancy (Fig. 3a: center panel) with no added phonons (brown curve) shows several Rabi oscillations at frequency $\Omega$, which is dominated by the detuning. Obvious changes in the dynamics are resolved when the SAW cavity is occupied such that $g/2\pi$=$770$ MHz (lowest red curve). These differences become more pronounced as $g$ increases, where the QD occupancy is estimated to reach 0.075 for $g/2\pi$=$2.4$ GHz at 800 ps. The data reveal that there are several pulse durations where phonon-assisted excitation (red curves) dominates substantially over the bare QD excitation (brown curve) — namely around 300 ps and 600 ps — largely governed by times when the bare QD occupancy swings down to approximately zero. These measured enhancements $c_{g}(t)$ reach between 12 and 45 at 600 ps, depending on the SAW cavity occupancy (Fig. 3a; bottom panel). To illustrate the importance of the pulse shape on the measured enhancements, we perform a similar set of measurements using an (approximately) square optical pulse (Fig. 3b). Although obvious changes in the dynamics are again observed with $g/2\pi$ as small as 340 MHz (bottom panel; red curve), the measured enhancements are considerably smaller than in the “exponential” pulse case, remaining near unity in all cases presented here.
The benefit of a gradually decreasing pulse intensity — as compared to a square pulse for example — becomes clear upon inspection of the bare QD dynamics. Numerical calculations of QD dynamics are presented in Figs. 3c-d, roughly corresponding to the two experimental cases shown in Figs. 3a-b (Methods). For a pulse with constant intensity (Fig. 3d; brown curve), the bare QD occupancy oscillates around and approaches a constant central value as radiative relaxation (or dephasing in general) proceeds. Consequently, successive minima in the bare QD’s occupancy grow increasingly large as the pulse continues. This behavior is also observed in our experiments (Fig. 3b). Enhancements remain smaller than 10 throughout the pulse duration. If the pulse intensity is instead gradually decreased soon after the onset, minima in the bare QD’s occupancy can grow increasingly close to zero over the pulse duration (Fig. 3c; brown curve). Enhancements thus increase over the course of the QD’s lifetime or longer, approaching maximum values near 1000 for $g/2\pi$$\gtrsim$1 GHz at long times. Calculations presented in Fig. 3 assume radiative relaxation and no additional dephasing (Methods); the effect of additional dephasing is discussed in the next section.
In our measurements, successive detection events sample different mechanical phases $\phi$, and so the occupancies shown in Fig. 3a,b represent the dynamics averaged over all such phases. Calculations shown in Figs. 3c,d are thus averaged over this phase. While $\phi$ does affect the specific trajectory, it has only a weak effect on the shape and magnitude of $c_{g}(t)$ which is our primary concern here (Supplementary Information). This scenario would in fact be relevant for, e.g., single-mode quantum microwave-to-optical transduction where the relevant phonon Fock states do not carry a well-defined phase. For applications where the specific trajectory should be well controlled, for example in the above-mentioned “swing-up” scheme, one can synchronize the mechanical drive phase with the onset of the optical pump [50].
IV Comparing pulsed and CW pumping schemes
Fig. 4 summarizes our experimental (open markers) and numerical (solid/dashed curves) results, where we plot maximal enhancements as a function of $g$. We theoretically consider three idealized and simple pulse shapes — square, “exponential”, and triangle, as illustrated in the top panels — and TLSs with both radiative relaxation only (solid curves) and additional pure dephasing (dashed curves); see Methods. “Exponential” and “square” correspond approximately to our two experimental cases, “exponential” referring to the exponentially decaying tail. With optical pulse lengths comparable to the TLS lifetime of $\approx$1 ns, triangular and “exponential” pulse shapes provide very similar performance, and both provide approximately ten-fold greater maximal enhancement than square pulses. The importance of both (1) a gradually decreasing pulse intensity and (2) lifetime-limited TLS dephasing are clear. Additional dephasing over timescales comparable to the QD lifetime — and thus the pulse duration — rapidly reduces maximal enhancements. Our experimental results evidence some form of dephasing in addition to radiative decay. Nonetheless, the experimentally obtained improvement between pulse shapes corresponds well to the ten-fold improvement seen in calculations. This suggests that our results are limited by slow spectral fluctuations rather than pure dephasing at the timescale of the QD’s lifetime. That is, our experimental occupancies are slightly “smeared” as the detuning effectively varies between successive detection events. This will be largely alleviated by investigating QDs with narrower measured linewidths.
[…]
V Discussion
We experimentally demonstrated how coherent optical control methods available with QD optomechanical systems can be exploited to enhance phonon-assisted scattering in applications such as microwave-to-optical transduction. Specifically, time-correlated single-photon counting measurements reveal excitation dynamics in a semiconductor QD that depend on the presence of microwave-frequency phonons. Pulse durations exist where background scattering from direct excitation is greatly reduced, improving the relative rate from phonon-assisted scattering. This signal-to-background ratio depends on the basic features of the optical pulse, including amplitude, detuning, shape, and phase. We explored three very simple distinct pulse shapes and demonstrated that pulses with a gradually decreasing intensity are beneficial. More complicated pulse shapes involving frequency, amplitude and phase shaping may be developed to further improve the performance.
More generally, this work shows that with the QD, new parameters are available for interacting with the mechanical system. For example, in addition to pump intensity, the performance now depends on details of the varying pump — including the amplitude and phase (not explored in this work) — at nanosecond timescales. In fact, with this time-domain control, there is no fundamental reason for matching the detuning to the mechanical mode frequency as is needed with bosonic resonators [2]. Using additional detuned pumps may offer another level of control, similar to previously demonstrated “swing-up” schemes [21, 32, 50]. Since the QD’s lifetime sets the timescale of state preparation and readout, average optical pump powers used for these protocols are at the nanowatt scale and this optical power is not stored in a resonator, reducing absorption-induced heating which is a major concern in optomechanical crystals [64, 13, 18].
Comparing our experimental results to numerical calculations suggests that our measured QD experiences slow spectral fluctuations that decrease the experimentally obtained enhancements. This is likely reflected in our measured QD linewidth of 800 MHz, but additional characterization would be needed to disentangle pure dephasing at the timescale of the QD’s lifetime from slow (sub-MHz) spectral wandering [60]. Modern optimized charge-controlled QDs regularly exhibit near-lifetime-limited linewidths [59, 57, 65], and we thus anticipate that performance comparable to our idealized numerical calculations is possible. Overall improved performance is anticipated from p-i-n charge-controlled QDs which offer greater charge and frequency stability, greater optical pumping efficiency, higher photon collection efficiency, and reduced pump background. Although our numerical calculations do not take into account the finite pump bandwidth, this is likely to play only a small role in experiments since the ideal pulse durations are longer than 1 ns, corresponding to transform-limited bandwidths of $\approx$350 MHz which are small compared to our detuning of $\approx$3.5 GHz. Larger issues likely arise from small polarization drifts in the experimental system which may change the field amplitude and temporal pulse shape that the QD experiences over the course of the 30-60 second integration times used in our experiments. Future efforts should involve optimizing QD performance in application-relevant mechanical resonators and exploring the broad parameter space of optical pumping schemes for improving the enhancements described here.
VI Acknowledgements
We thank Emanuel Knill for informative discussions and intellectual support related to this work.
VII Author Contributions
R.A.D., A.K., and K.L.S. conceived and designed the experiments. R.A.D. performed the measurements and analyzed the data. All authors contributed materials and analysis tools. R.A.D. and K.L.S. wrote the paper with input from all other authors.
VIII Methods
VIII.1 System Hamiltonian
The Hamiltonian describing our system, in a frame rotating with the optical pump at frequency $\omega_{\text{pump}}$, is given by
$\frac{\tilde{\hat{H}}(t)}{\hbar}=\frac{1}{2}\left[-\Delta+g\cos(\omega_{\text{SAW}}t+\phi)\right]\hat{\sigma}_{z}+\frac{\Omega_{0}(t)}{2}\hat{\sigma}_{x}$ (1)
where $\Delta$, $\Omega_{0}$ and $g$ describe the pump-QD detuning, the resonant Rabi rate, and QD-SAW coupling, respectively. $\Omega_{0}$ is taken to be a real-valued function with a time dependence that describes the optical pulse envelope. We numerically calculate the dynamics of the system, described by Eqn. 1, using the Lindblad master equation with the open-source Python library QuTiP. Radiative decay is included through the Lindblad operator $\hat{\sigma}_{-}=(\hat{\sigma}_{x}+i\hat{\sigma}_{y})/2$ with rate $\gamma_{\text{QD}}$. Additional pure dephasing is included through the Lindblad operator $\hat{\sigma}_{z}$ with rate $\gamma_{z}$. The rates $\gamma_{\text{QD}}$ and $\gamma_{z}$ are specified in the manuscript where relevant. Further details are provided in the Supplementary Information.
VIII.2 Experimental Setup
All of our QD optical measurements were performed using a 5 K optical cryostat with an internal cryogenic 0.81 NA objective. The resonant pump beam is shaped outside the cryostat using single-mode optical fiber components and free-space optical components. The resonant pump beam was a tunable external-cavity diode laser with a 10 kHz linewidth. Pump beams illuminate the QD through the objective. QD luminescence was collected by the same objective and isolated from the pump predominantly by spectral filtering (for measurements involving a nonresonant pump) or polarization filtering (for measurements involving a resonant pump) with at least 106 pump rejection. Luminescence was collected into single-mode optical fiber and various frequency- and time-domain measurements were performed using procedures detailed below. The QD bias voltage was held constant for all measurements shown in the manuscript. The QD emission wavelength (and therefore the resonant pump wavelength) used in this work was 975.934 nm.
VIII.3 Continuous-wave Resonance Fluorescence Spectra
Continuous-wave (CW) resonance fluorescence (RF) spectra (e.g. Fig. 2) were recorded by slowly sweeping the frequency of a tightly focused resonant pump beam over the QD’s exciton resonance while holding the QD bias constant. Luminescence counts at each frequency were recorded using a superconducting nanowire single photon detector (SNSPD) mounted in a dilution refrigerator. A pump power of approximately 2 nW at the sample was used, corresponding to approximately 1/10 the measured saturation level of the QD. For manuscript Fig. 2d (upper panel), such spectra were recorded without SAW driving and with SAW driving using an external coherent microwave source at approximately 3.518 GHz and external microwave driving powers between -50 dBm and -35 dBm. External microwave driving powers were converted to coupling rates $g$ by a series of calibration data where the modulation index $\chi$ was fit for each microwave driving power and $g$ was derived by the relationship $g$=$\omega_{\text{SAW}}\chi$. The calibration data are shown in the Supplementary Information.
For manuscript Fig. 2d (lower panel), the “resonant” pump was held at a constant frequency detuning of $\Delta$=$-\omega_{\text{SAW}}$. A pump power of approximately 5 nW was used. The frequency spectrum of scattered photons was measured by passing the collection port through a tunable Fabry-Perot etalon filter with a 600 MHz linewidth before being counted on an SNSPD. In manuscript Fig. 2d (lower panel), such spectra were again measured without SAW driving and with SAW driving at 3.518 GHz and the same powers as in manuscript Fig. 2d (upper panel).
VIII.4 Pulsed Optical Measurements and Time-domain QD Dynamics
The time evolution of the QD dynamics was measured using standard time-correlated single-photon counting (TCSPC) techniques with a two-channel picosecond event timer and TCSPC module. For “exponential” decaying pulses, optical pulses were generated by passing a CW “resonant” optical pump (same as described above) through two cascaded electro-optic modulators (EOMs). The EOMs were driven by amplified pulses from an approximately 100-ps square-wave electronic pulse generator (EPG). The EPG was externally triggered using an arbitrary waveform generator (AWG) at 38 MHz. The rise time, fall time and width of the output pulses were measured to be approximately 15 ps, 15 ps, and 130 ps respectively. The on/off contrast was measured to be better than $10^{4}$. The frequency bandwidth of the output pulses was measured to be 7 GHz. To decrease the bandwidth (and increase the pulse duration), the pulses were transmitted through a 600 MHz tunable Fabry-Perot etalon which was tuned for maximum transmission at the pulse’s center frequency. For square pulses, the EOMs were directly driven using a high-speed AWG (64 GS/sec) with approximately 30 ps rise time and pulses of arbitrary length. The pulse center frequency was detuned by $\Delta$ with respect to the QD’s resonance frequency (detuning specified in each measurement). Measured pulse profiles shown in the manuscript were from the reflected signal and may contain system reflections that don’t illuminate the QD. That is, the pulse profile experienced by the QD may differ slightly from what we measured. This may partially account for the onset of the QD’s free evolution around 800 ps in Fig. 3a although the measured pulse continues beyond this time.
The time distribution of the QD luminescence was measured by again detecting the luminescence with an SNSPD and time-correlating the counts on the TCSPC module with respect to a trigger signal from the AWG. We typically used 30-60 second total integration times, over which we collected typically several million total photons. This type of measurement was performed with a wide range of average optical input powers (between approximately 0.25 nW to 1500 nW). Pump powers were converted to resonant Rabi rates by looking at the resulting time-domain oscillations.
The time-resolved QD dynamics shown in the manuscript are presented after subtracting the experimentally measured pump contribution (i.e. unrejected pump background). Count rates were converted to approximate excited state occupancies using a series of power-dependent calibration data where observed Rabi oscillations were compared to numerical calculations. This subtraction process and scaling process are further illustrated in the Supplementary Information.
Supplementary Information
1 System Hamiltonian
Our system comprises a two-level system (TLS) with frequency $\omega_{\text{QD}}$ driven at optical frequency $\omega_{\text{pump}}$ with a coupling (Rabi) rate $\Omega_{0}$. The TLS frequency is modulated at microwave frequency $\omega_{\text{SAW}}$ with a modulation depth $\chi$=$g/\omega_{\text{QD}}$. The Hamiltonian $\hat{H}$ is:
$\frac{\hat{H}(t)}{\hbar}=\frac{1}{2}\omega_{\text{QD}}\hat{\sigma}_{z}+\frac{1}{2}g\cos(\omega_{\text{SAW}}t+\phi)\hat{\sigma}_{z}+\Omega_{0}\cos(\omega_{\text{pump}}t)\hat{\sigma}_{x}$ (2)
[…]
$\frac{\tilde{\hat{H}}(t)}{\hbar}=\frac{1}{2}\left[-\Delta+g\cos(\omega_{\text{SAW}}t+\phi)\right]\hat{\sigma}_{z}+\frac{\Omega_{0}(t)}{2}\hat{\sigma}_{x}$ (3)
We have allowed $\Omega_{0}$ to adopt a time dependence which describes the optical pulse envelope. Formally, Eqn. 2 assumes that the time variation of $\Omega_{0}(t)$ is slow enough that the resulting spectral distribution remains small compared to the detuning $\Delta$ so that the optical drive can be approximated as monochromatic. Experimentally, we enforce this using long pulses and/or spectral filters. We additionally assume that $\Omega_{0}(t)$ is a real-valued function. In general, pulse shaping will cause the pulse amplitude to become complex which may affect the specific TLS trajectories.
We numerically calculate the dynamics of the system, described by Eqn. 2, using the Lindblad master equation for the density matrix $\rho(t)$:
$\dot{\rho}(t)=-\frac{i}{\hbar}[\tilde{\hat{H}}(t),\rho(t)]+\frac{1}{2}\sum_{n}[2\hat{C}_{n}\rho(t)\hat{C}_{n}^{\dagger}-\rho(t)\hat{C}_{n}^{\dagger}\hat{C}_{n}-\hat{C}_{n}^{\dagger}\hat{C}_{n}\rho(t)]$ (4)
with collapse operators $\hat{C}_{n}$=$\sqrt{\gamma_{n}}\hat{\sigma}_{n}$ and rates $\gamma_{n}$. Radiative decay is included through the Lindblad operator $\hat{\sigma}_{-}=(\hat{\sigma}_{x}+i\hat{\sigma}_{y})/2$ with rate $\gamma_{\text{QD}}$. Additional pure dephasing is included through the Lindblad operator $\hat{\sigma}_{z}$ with rate $\gamma_{z}$. Note that “slow” spectral fluctuations — those occurring over timescales significantly longer than the QD’s lifetime — affect the measured TLS trajectories in a qualitatively different way than pure dephasing. We did not account for such spectral fluctuations in our calculations. Calculations are performed using the open-source Python library QuTiP.
The mechanical phase $\phi$ affects the specific trajectory around the Bloch sphere. As described in the manuscript, our measurements sample many instances of $\phi$ and so the presented occupancy dynamics represent averages over all $\phi$. Calculations shown in manuscript Figs. 3c-d account for this by averaging calculated trajectories over eight uniformly spaced values of $\phi$ ranging from 0 to $2\pi$. Figs. S1a,b illustrate how the trajectory changes with $\phi$ (panel a), and also the result of averaging over these trajectories (panel b). Figs. S1c,d show that this averaging has only weak effects on the occupancy ratio $c_{g}(t)$ although the specific dynamics are changed.
Figure S1: (a) Calculated TLS occupancies for five different values of $\phi$ (solid curves). The $g=0$ trajectory (which does not depend on $\phi$) is also shown for reference (dotted curve) (b) Calculated TLS occupancies for $\phi$=0 (blue curve) and for $\phi$ averaged over eight uniformly-spaced values between 0 and $2\pi$ (orange curve). All calculations in panels a,b were performed for an idealized square pulse with the following parameters: $g/2\pi$=1 GHz, $\Omega_{0}$=1 GHz. (c,d) Occupancies and occupancy ratios ($c_{g}(t)$) calculated for the same set of parameters as in manuscript Fig. 3d for (c) $\phi$=0 and for (d) $\phi$ averaged over eight values between 0 and $2\pi$.
2 Experimental Setup
All of our QD optical measurements were performed using a 5 K optical cryostat with an internal cryogenic 0.81 NA objective. Nonresonant and resonant pump beams were shaped outside the cryostat using single-mode optical fiber components and free-space optical components. The resonant pump beam was a tunable external-cavity diode laser with an approximately 10 kHz linewidth at 970 nm. The nonresonant pump was a 632 nm diode laser. Pump beams illuminate the QD through the objective. QD luminescence was collected by the same objective and isolated from the pump predominantly by spectral filtering (for measurements involving a nonresonant pump) or polarization filtering (for measurements involving a resonant pump) with at least 106 pump rejection. Luminescence was collected into single-mode optical fiber and various frequency- and time-domain measurements were performed using procedures detailed below. The QD bias voltage was held constant for all measurements except for Fig. 2c of the manuscript.
2.1 Bias-dependent QD Photoluminescence Spectra
QDs were illuminated with nonresonant pump light. Photoluminescence (PL) was collected into a single-mode fiber and reflected pump light was spectrally rejected. Pump powers were kept well below saturation levels. PL was sent to a spectrometer and recorded with a CCD camera. The QD bias was varied in discrete steps while repeated integrations were performed. Spectra were stitched together in post-processing. The bias-dependent PL spectrum from the QD interrogated in the manuscript is shown in Fig. S2.
Figure S2: QD PL spectrum with varying Schottky bias voltage. Two to three QDs contribute to this spectrum. The QD investigated in the manuscript is indicated by a white arrow.
2.2 Continuous-wave Resonance Fluorescence Spectra
Continuous-wave (CW) resonance fluorescence (RF) spectra (e.g. manuscript Fig. 2a) were recorded by slowly sweeping the frequency of a tightly focused resonant pump beam over the QD’s exciton resonance while holding the QD bias constant. Luminescence counts at each frequency were recorded using a superconducting nanowire single photon detector (SNSPD) mounted in a dilution refrigerator. A pump power of approximately 2 nW at the sample was used, corresponding to approximately 1/10 the measured saturation level of the QD. For manuscript Fig. 2d (upper panel), such spectra were recorded without SAW driving and with SAW driving using an external coherent microwave source at approximately 3.518 GHz and external microwave driving powers between -50 dBm and -35 dBm. External microwave driving powers were converted to coupling rates $g$ by a series of calibration data where the modulation index $\chi$ was fit for each microwave driving power and $g$ was derived by the relationship $g$=$\omega_{\text{SAW}}\chi$. The calibration data are shown in Fig. S3.
For manuscript Fig. 2d (lower panel), the “resonant” pump was held at a constant frequency detuning of $\Delta$=$-\omega_{\text{SAW}}$. A pump power of approximately 5 nW was used. The frequency spectrum of scattered photons was measured by passing the collection port through a tunable Fabry-Perot etalon filter with a 600 MHz linewidth before being counted on an SNSPD. In manuscript Fig. 2d (lower panel), such spectra were again measured without SAW driving and with SAW driving at 3.518 GHz and the same powers as in manuscript Fig. 2d (upper panel).
Figure S3: Calibration between microwave driving power $P_{\mu}$ and coupling rate $g$. Orange markers: data. Blue curve: fit assuming $g\propto\sqrt{P_{\mu}}$.
2.3 Pulsed Optical Measurements and Time-domain QD Dynamics
The time evolution of the QD dynamics was measured using standard time-correlated single-photon counting (TCSPC) techniques with a two-channel picosecond event timer and TCSPC module. For “exponential” decaying pulses, optical pulses were generated by passing a CW “resonant” optical pump (same as described above) through two cascaded electro-optic modulators (EOMs). The EOMs were driven by amplified pulses from an approximately 100-ps square-wave electronic pulse generator (EPG). The EPG was externally triggered using an arbitrary waveform generator (AWG). The rise-time, fall-time and width of the output pulses were measured to be approximately 15 ps, 15 ps, and 130 ps respectively. The on/off contrast was measured to be better than $10^{4}$. The frequency bandwidth of the output pulses was measured to be 7 GHz. To decrease the bandwidth (and increase the pulse duration), the pulses were transmitted through a 600 MHz tunable Fabry-Perot etalon which was tuned for maximum transmission at the pulse’s center frequency. For square pulses, the EOMs were directly driven using a high-speed AWG (64 GS/sec) with approximately 30 ps rise time and pulses of arbitrary length. The pulse center frequency was detuned by $\Delta$ with respect to the QD’s resonance frequency (detuning specified in each measurement). Measured pulse profiles shown in the manuscript may differ slightly from the pulse experienced by the QD due to small system reflections outside the cryostat. The time distribution of the QD luminescence was measured by again detecting the luminescence with an SNSPD and time-correlating the counts on the TCSPC module with respect to a trigger signal from the AWG. We typically used 30-60 second total integration times, over which we collected typically several million total photons. This type of measurement was performed with a wide range of average optical input powers (between approximately 0.25 nW to 1500 nW). Pump powers were converted to resonant Rabi rates by looking at the resulting time-domain oscillations.
The time-resolved QD dynamics shown in the manuscript are presented after subtracting the experimentally measured pump contribution. This subtraction process is illustrated in Fig. S4, where black curves represent the directly measured data including pump and QD luminescence contributions, gray regions represent the measured pump contribution, and blue curves represent the difference which we associated with the QD contribution.
Figure S4: A series of raw TCSPC measurements showing measured pump background (gray filled region), QD signal (black curve) and pump-subtracted QD signal (blue curve). Each panel includes a different set of experimental parameters specified above each panel.
3 Estimating QD Occupancies from Experimental Time-Domain Data
QD occupancies were derived by comparing count rates from time-resolved QD luminescence data to a reference count rate. The reference count rate was derived by performing a set of time-resolved QD luminescence measurements with 130 ps optical pulses and widely varying optical pump powers and analyzing the resulting Rabi oscillations in the signal. Experimental Rabi oscillations were compared to numerical calculations of the QD occupancy under a 130-ps square pulse. The calculation includes dephasing assuming that the QD dephasing is purely radiative. (The measured 1 ns radiative lifetime corresponds to a 321 MHz decay rate.) This calibration is illustrated in Fig. S5. Note that experimental occupancies will be overestimated if the QD experiences additional dephasing. Nonetheless, this scaling is irrelevant for the enhancements presented in the manuscript.
Figure S5: Left panel: Raw TCSPC data acquired with $\Delta=0$ and a 130 ps pulse duration. “Maximum counts” refers to the counts recorded in a 16-ps time bin at 140 ps, i.e., immediately after the pump pulse. Right panel: Counts recorded per second within a 16-ps time bin at 140 ps as a function of average optical power (open black square markers). Single-time-bin counts are converted to an expected QD occupancy $\langle\rho_{ee}\rangle$ by comparing the resulting Rabi oscillations to theoretical calculations including lifetime-limited dephasing at rate $\gamma_{\text{QD}}/2\pi$=320 MHz (solid black curve).